OpenStudy (anonymous):

William got $150 for making all A’s on his report card. He wants to spend it on CDs and DVDs. The local video store is having a sale; all CDs are $10 and all DVDs are $15. Which of the following is a possible combination of CDs and DVDs? A. 2 CDs and 9 DVDs B. 7 CDs and 5 DVDs C. 8 CDs and 6 DVDs D. 11 CDs and 3 DVDs

OpenStudy (texaschic101):

10c + 15d = 150 do the same thing like the last one we did.....sub in your answer choices and see which one fits

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

your smart....you can do this

OpenStudy (anonymous):

its B 70+75=145

OpenStudy (texaschic101):

correct ....I told you you could do this :)

OpenStudy (anonymous):

A:20+135=155 B:70+75=145 C:80+90=170 D:110+45=155

OpenStudy (anonymous):

Thanks